# Policy-Driven World Model Adaptation for Robust Offline Model-based Reinforcement Learning

### v2

@@ Appendix E Discussion on the Size of the Function Class
Next, we show that for tabular MDPs considered in Theorem 2, $J(\theta^{*},\phi^{*})-J(\hat{\theta},\phi^{*})\leq\frac{\sqrt{C_{\phi^{*},\theta^{*}}}}{(1-\gamma)^{2}}\sqrt{4\epsilon+c^{\prime}\left(\sqrt{\frac{\log(2N_{c}/\delta)}{N}}+\frac{\log(2N_{c}/\delta)}{N}\right)}$, with probability at least $1-\delta$, where $c^{\prime}$ is a constant and $N_{c}$ is upper bounded by ${KN+K-1\choose K-1}^{|\mathcal{S}||\mathcal{A}|}$.
@@
For a fixed $(s,a)$, $P_{\phi}(\cdot|s,a)\in\mathbb{R}^{K}$ is a categorical distribution. According to Lemma 5, the $\frac{1}{N}$-covering number for $\{P_{\phi}(\cdot|s,a)\}$ in terms of the L1-norm is upper bounded by $KN+K-1\choose K-1$}. Then, we can construct a set $\{P_{1},\cdots,P_{N_{c}}\}$, such that $\forall P_{\phi}$, there exists $P_{i}$ such that $||P_{i}(\cdot|s,a)-P_{\phi}(\cdot|s,a)||_{1}\leq\frac{1}{N}$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$, and we have $N_{c}\leq{KN+K-1\choose K-1}^{|\mathcal{S}||\mathcal{A}|}$.
@@
For MDPs with high-dimensional continuous state and action spaces, it is common practice to learn deep neural network-based world models to ensure that the function space $\mathcal{M}$ adequately represents the true environment MDP and its maximum likelihood estimator [DBLP:journals/corr/abs-2112-14523]. An upper bound on $|\Phi|$ can be similarly derived using the approach outlined above, leveraging bounds on the covering number for deep neural networks [DBLP:conf/icml/Shen24]. However, as this falls outside the primary focus of this paper – developing a practical algorithm for robust offline MBRL – we leave it as an important direction for future work.
@@ Appendix F Derivation of Equation (9)
As in Appendix A and [DBLP:conf/icml/FiezCR20], by relacing $(\theta,\phi^{*},\lambda^{*})$ with $(\theta^{k},\phi^{k},\lambda^{k})$, we obtain the gradient update for $\theta$ at iteration $k$, corresponding to the first line of Eq. (9).
Appendix G Proof of Theorem 4
@@
Finally, based on Theorem 4 and the definition of $\mathcal{L}(\theta,\phi,\lambda)$ (in Section 4), we can get the following derivatives, which are used in Eqs. (7) - (9):
@@ Appendix I Analysis of Time Complexity
Algorithm 1 differs from typical offline MBRL algorithms, such as [DBLP:conf/icml/SunZJLYY23, DBLP:journals/corr/abs-2410-11234], in several key aspects. First, they learn an ensemble of world models through supervised learning before the MBRL process, as in Line 2 of Algorithm 1. However, unlike our approach, the world models are not updated alongside the policy during MBRL. Second, they introduce an additional penalty term to the predicted reward from $P_{\bar{\phi}}$. This penalty is large when there is high variance among the predictions from different ensemble members, discouraging the agent from visiting uncertain regions in the state-action space. Third, at each training iteration, they randomly sample a set of states from $\mathcal{D}_{\mu}$ as starting points for imaginary rollouts. During these rollouts, the policy $\pi_{\theta}$ interacts with $P_{\bar{\phi}}$ for a short horizon of length $l$, collecting transitions into a replay buffer for RL. The use of short-horizon rollouts helps mitigate compounding errors caused by inaccurate world model predictions. Finally, they train $\pi_{\theta}$ for multiple epochs using transitions stored in the buffer with Soft Actor-Critic (SAC) [DBLP:conf/icml/HaarnojaZAL18, chen2023multiagentdeepcoveringskill], an off-policy RL algorithm.
@@
Given that the underlying world model is continuously updated, maintaining a large replay buffer and computing policy gradients based on outdated, off-policy trajectories, as done in SAC, becomes infeasible. To improve sample efficiency, instead of updating the policy and world model only once per learning iteration, as done in Lines 6 and 7 of Algorithm 1, we can update them for multiple epochs using the same batch of on-policy samples. Following PPO [DBLP:journals/corr/SchulmanWDRK17, DBLP:journals/corr/abs-2405-16386], we apply gradient masks to both the policy and world model to regulate parameter updates based on outdated data. Specifically, the policy gradient mask for $(s_{t},a_{t})$ within a trajectory $\tau$ is defined as:
@@
Scalar States
$\beta_{N}$, Internal Inductance, Line Averaged Density,
Loop Voltage, Stored Energy
Profile States
Electron Density, Electron Temperature, Pressure,
Safety Factor, Ion Temperature, Ion Rotation
@@
Elongation, Top Triangularity, Bottom Triangularity, Minor Radius,
Radius and Vertical Locations of the Plasma Center
Direct Actuators
Power Injected, Torque Injected, Total Deuterium Gas Injection,
@@
Total ECH Power, Magnitude and Sign of the Toroidal Magnetic Field
@@
Nuclear fusion is a promising energy source to meet the world’s growing demand. It involves fusing the nuclei of two light atoms, such as hydrogen, to form a heavier nucleus, typically helium, releasing energy in the process. The primary challenge of fusion is confining a plasma, i.e., an ionized gas of hydrogen isotopes, while heating it and increasing its pressure to initiate and sustain fusion reactions. The tokamak is one of the most promising confinement devices. It uses magnetic fields acting on hydrogen atoms that have been ionized (given a charge) so that the magnetic fields can exert a force on the moving particles [1512794].
The authors of [DBLP:journals/corr/abs-2404-12416] trained an ensemble of deep recurrent probabilistic neural networks as a surrogate dynamics model for the DIII-D tokamak, a device located in San Diego, California, and operated by General Atomics, using a large dataset of operational data from that device. A typical shot (i.e., episode) on DIII-D lasts around 6-8 seconds, consisting of a one-second ramp-up phase, a multi-second flat-top phase, and a one-second ramp-down phase. The DIII-D also features several real-time and post-shot diagnostics that measure the magnetic equilibrium and plasma parameters with high temporal resolution. The authors demonstrate that the learned model predicts these measurements for entire shots with remarkable accuracy. Thus, we use this model as a "ground truth" simulator for tokamak control tasks. Specifically, we generate a dataset of 111305 transitions for offline RL by replaying actuator sequences from real DIII-D operations through the ensemble of dynamics models. Policy evaluation is conducted using a noisy version of this data-driven simulator to assess deployment robustness.
The state and action spaces for the tokamak control tasks are outlined in Table 3. For detailed physical explanations of them, please refer to [abbate2021data, DBLP:conf/l4dc/CharABBCCEMRKS23, ariola2008magnetic]. The state space consists of five scalar values and six profiles which are discretized measurements of physical quantities along the minor radius of the toroid. After applying principal component analysis [mackiewicz1993principal], the pressure profile is reduced to two dimensions, while the other profiles are reduced to four dimensions each. In total, the state space comprises 27 dimensions. The action space includes direct control actuators for neutral beam power, torque, gas, ECH power, current, and magnetic field, as well as target values for plasma density and plasma shape, which are managed through a lower-level control module. Altogether, the action space consists of 14 dimensions. Following guidance from fusion experts, we select a subset of the state space: $\beta_{N}$ and all profile states, as policy inputs, spanning 23 dimensions. The policy is trained to control five direct actuators: power, torque, total ECH power, the magnitude and sign of the toroidal magnetic field. For the remaining actuators listed in Table 3, we replay the corresponding sequences from recorded DIII-D operations.

### v3

@@ Appendix E Discussion on the Size of the Function Class
Next, we show that for tabular MDPs considered in Theorem 3.2, $J(\theta^{*},\phi^{*})-J(\hat{\theta},\phi^{*})\leq\frac{\sqrt{C_{\phi^{*},\theta^{*}}}}{(1-\gamma)^{2}}\sqrt{4\epsilon+c^{\prime}\left(\sqrt{\frac{\log(2N_{c}/\delta)}{N}}+\frac{\log(2N_{c}/\delta)}{N}\right)}$, with probability at least $1-\delta$, where $c^{\prime}$ is a constant and $N_{c}$ is upper bounded by ${KN+K-1\choose K-1}^{|\mathcal{S}||\mathcal{A}|}$.
@@
For a fixed $(s,a)$, $P_{\phi}(\cdot|s,a)\in\mathbb{R}^{K}$ is a categorical distribution. According to Lemma E.1, the $\frac{1}{N}$-covering number for $\{P_{\phi}(\cdot|s,a)\}$ in terms of the L1-norm is upper bounded by $KN+K-1\choose K-1$}. Then, we can construct a set $\{P_{1},\cdots,P_{N_{c}}\}$, such that $\forall P_{\phi}$, there exists $P_{i}$ such that $||P_{i}(\cdot|s,a)-P_{\phi}(\cdot|s,a)||_{1}\leq\frac{1}{N}$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$, and we have $N_{c}\leq{KN+K-1\choose K-1}^{|\mathcal{S}||\mathcal{A}|}$.
@@
For MDPs with high-dimensional continuous state and action spaces, it is common practice to learn deep neural network-based world models to ensure that the function space $\mathcal{M}$ adequately represents the true environment MDP and its maximum likelihood estimator (Beneventano et al., 2021). An upper bound on $|\Phi|$ can be similarly derived using the approach outlined above, leveraging bounds on the covering number for deep neural networks (Shen, 2024). However, as this falls outside the primary focus of this paper – developing a practical algorithm for robust offline MBRL – we leave it as an important direction for future work.
@@ Appendix F Derivation of Equation (9)
As in Appendix A and (Fiez et al., 2020), by relacing $(\theta,\phi^{*},\lambda^{*})$ with $(\theta^{k},\phi^{k},\lambda^{k})$, we obtain the gradient update for $\theta$ at iteration $k$, corresponding to the first line of Eq. (9).
Appendix G Proof of Theorem 5.1
@@
Finally, based on Theorem 5.1 and the definition of $\mathcal{L}(\theta,\phi,\lambda)$ (in Section 4), we can get the following derivatives, which are used in Eqs. (7) - (9):
@@ Appendix I Analysis of Time Complexity
Algorithm 1 differs from typical offline MBRL algorithms, such as (Sun et al., 2023; Chen et al., 2024a), in several key aspects. First, they learn an ensemble of world models through supervised learning before the MBRL process, as in Line 2 of Algorithm 1. However, unlike our approach, the world models are not updated alongside the policy during MBRL. Second, they introduce an additional penalty term to the predicted reward from $P_{\bar{\phi}}$. This penalty is large when there is high variance among the predictions from different ensemble members, discouraging the agent from visiting uncertain regions in the state-action space. Third, at each training iteration, they randomly sample a set of states from $\mathcal{D}_{\mu}$ as starting points for imaginary rollouts. During these rollouts, the policy $\pi_{\theta}$ interacts with $P_{\bar{\phi}}$ for a short horizon of length $l$, collecting transitions into a replay buffer for RL. The use of short-horizon rollouts helps mitigate compounding errors caused by inaccurate world model predictions. Finally, they train $\pi_{\theta}$ for multiple epochs using transitions stored in the buffer with Soft Actor-Critic (SAC) (Haarnoja et al., 2018; Chen et al., 2023b), an off-policy RL algorithm.
@@
Given that the underlying world model is continuously updated, maintaining a large replay buffer and computing policy gradients based on outdated, off-policy trajectories, as done in SAC, becomes infeasible. To improve sample efficiency, instead of updating the policy and world model only once per learning iteration, as done in Lines 6 and 7 of Algorithm 1, we can update them for multiple epochs using the same batch of on-policy samples. Following PPO (Schulman et al., 2017; Chen et al., 2024b), we apply gradient masks to both the policy and world model to regulate parameter updates based on outdated data. Specifically, the policy gradient mask for $(s_{t},a_{t})$ within a trajectory $\tau$ is defined as:
@@
Scalar States $\beta_{N}$, Internal Inductance, Line Averaged Density, Loop Voltage, Stored Energy
Profile States Electron Density, Electron Temperature, Pressure, Safety Factor, Ion Temperature, Ion Rotation
@@
Shape Variables Elongation, Top Triangularity, Bottom Triangularity, Minor Radius, Radius and Vertical Locations of the Plasma Center
@@
Direct Actuators Power Injected, Torque Injected, Total Deuterium Gas Injection, Total ECH Power, Magnitude and Sign of the Toroidal Magnetic Field
Nuclear fusion is a promising energy source to meet the world’s growing demand. It involves fusing the nuclei of two light atoms, such as hydrogen, to form a heavier nucleus, typically helium, releasing energy in the process. The primary challenge of fusion is confining a plasma, i.e., an ionized gas of hydrogen isotopes, while heating it and increasing its pressure to initiate and sustain fusion reactions. The tokamak is one of the most promising confinement devices. It uses magnetic fields acting on hydrogen atoms that have been ionized (given a charge) so that the magnetic fields can exert a force on the moving particles (Pironti and Walker, 2005).
The authors of (Char et al., 2024) trained an ensemble of deep recurrent probabilistic neural networks as a surrogate dynamics model for the DIII-D tokamak, a device located in San Diego, California, and operated by General Atomics, using a large dataset of operational data from that device. A typical shot (i.e., episode) on DIII-D lasts around 6-8 seconds, consisting of a one-second ramp-up phase, a multi-second flat-top phase, and a one-second ramp-down phase. The DIII-D also features several real-time and post-shot diagnostics that measure the magnetic equilibrium and plasma parameters with high temporal resolution. The authors demonstrate that the learned model predicts these measurements for entire shots with remarkable accuracy. Thus, we use this model as a "ground truth" simulator for tokamak control tasks. Specifically, we generate a dataset of 111305 transitions for offline RL by replaying actuator sequences from real DIII-D operations through the ensemble of dynamics models. Policy evaluation is conducted using a noisy version of this data-driven simulator to assess deployment robustness.
The state and action spaces for the tokamak control tasks are outlined in Table 4. For detailed physical explanations of them, please refer to (Abbate et al., 2021; Char et al., 2023; Ariola et al., 2008). The state space consists of five scalar values and six profiles which are discretized measurements of physical quantities along the minor radius of the toroid. After applying principal component analysis (Maćkiewicz and Ratajczak, 1993), the pressure profile is reduced to two dimensions, while the other profiles are reduced to four dimensions each. In total, the state space comprises 27 dimensions. The action space includes direct control actuators for neutral beam power, torque, gas, ECH power, current, and magnetic field, as well as target values for plasma density and plasma shape, which are managed through a lower-level control module. Altogether, the action space consists of 14 dimensions. Following guidance from fusion experts, we select a subset of the state space: $\beta_{N}$ and all profile states, as policy inputs, spanning 23 dimensions. The policy is trained to control five direct actuators: power, torque, total ECH power, the magnitude and sign of the toroidal magnetic field. For the remaining actuators listed in Table 4, we replay the corresponding sequences from recorded DIII-D operations.
Appendix L Additional Ablation and Sensitivity Analysis
L.1 Sensitivity to Uncertainty Radius ($\epsilon$)
We evaluated ROMBRL on hopper-medium-replay under different noise injection levels ($\delta\in\{0.05,0.1,0.2\}$) with varying uncertainty radii $\epsilon\in\{1,5,10,20,50\}$. The results are plotted in Figure 4.
We observe a clear correlation between the optimal $\epsilon$ and the noise intensity. At low noise levels ($0.05$), a small radius is sufficient. However, as the testing noise increases to $0.2$, using a small $\epsilon$ (e.g., $\epsilon=1$) results in sub-optimal performance, as the constrained uncertainty set is too narrow to encompass the actual dynamics mismatch. Increasing $\epsilon$ (e.g., to $20$ or $50$) effectively recovers robustness. This validates the intuition that a larger uncertainty set is required to defend against larger environmental perturbations.
Figure 4: Sensitivity Analysis of Uncertainty Radius ($\epsilon$). Performance under varying testing noise levels. Larger noise levels require a larger $\epsilon$ to maintain robustness, while an overly small $\epsilon$ leads to performance degradation.
L.2 Verification of Two-Timescale Learning Rates
Our theoretical analysis in Section 4 suggests that the follower (world model) must converge significantly faster than the leader (policy) to ensure the leader observes a valid best response. To verify this, we conducted a grid search over policy learning rates $\eta_{\theta}$ and model learning rates $\eta_{\phi}$ on hopper-medium-replay under 5% noise injections. The results are visualized in Figure 5.
The experiment strictly aligns with our theoretical derivation. We observe a sharp performance collapse (score dropping to $\approx 3.3$) when the condition $\eta_{\theta}\geq\eta_{\phi}$ occurs (e.g., $\eta_{\theta}=10^{-2},\eta_{\phi}=10^{-2}$). High performance is consistently achieved only in the region where the two-timescale rule $\eta_{\phi}\gg\eta_{\theta}$ is satisfied (e.g., $\eta_{\phi}=10^{-2},\eta_{\theta}=10^{-4}$), ensuring stable convergence of the Stackelberg game.
Figure 5: Impact of Learning Rates. The heatmap shows average returns for different combinations of $\eta_{\theta}$ and $\eta_{\phi}$. Performance collapses when $\eta_{\theta}\geq\eta_{\phi}$, verifying the necessity of the two-timescale update rule.
Appendix M Experimental Setup Details
To ensure statistical validity and reproducibility, all reported results are the mean and standard deviation over 5 independent random seeds. We adhered to a principle of fair comparison for hyperparameter selection. For the D4RL MuJoCo tasks, all baselines were configured using the recommended hyperparameters provided in the official OfflineRL-Kit repository . For our method, ROMBRL, its base architectural hyperparameters are inherited from MOBILE, while its core robustness parameter, the uncertainty radius $\epsilon$, was kept fixed at 10 for all experiments without any task-specific tuning.
For the more challenging Tokamak Control tasks, no algorithm-specific hyperparameter tuning was performed for any method, including our own. Instead, to maintain a consistent and unbiased evaluation setting, we adopted the hyperparameters from a recognized difficult task within the MuJoCo suite, Walker2d-med-exp . This choice was motivated by its medium-expert data quality, which most closely resembles the dataset characteristics of the Tokamak tasks.
Table 5: Comparison of computational cost (wall-clock time per training epoch) and GPU memory usage, averaged over the -expert D4RL tasks. All experiments were conducted on a single NVIDIA Tesla V100 GPU.
Metric ROMBRL (ours) MOBILE RAMBO
Runtime (ms/epoch) 31.85 17.97 28.12
Memory Usage (MB) 1597 1586 1609
Appendix N Additional Experimental Results
In this section, we provide supplementary experimental results to address the valuable feedback from the reviewers. These include a comparative analysis of computational costs, an evaluation of our algorithm’s robustness under a higher noise intensity, and detailed training curves for all D4RL MuJoCo tasks.
N.1 Computational Cost Analysis
To address concerns regarding computational efficiency, we provide a comparison of the wall-clock runtime per training epoch and the peak GPU memory usage. The experiments were conducted on a single NVIDIA Tesla V100-SXM2-32GB GPU, with results averaged over the expert-level tasks for Hopper, Walker2d, and HalfCheetah. As shown in Table 5, the memory usage of ROMBRL is comparable to both MOBILE and RAMBO. While ROMBRL’s runtime is higher than that of MOBILE, it is comparable to RAMBO, another baseline specifically designed for robustness. This indicates that the computational overhead of our method is within a reasonable range for robust offline RL algorithms. Additionally, our approach requires no additional components during deployment, and this strategy performs comparably to all other baselines in real-world tasks.
N.2 Robustness under Higher Noise Levels
While our main experiments use a 5% noise level, which is a common assumption for sensor measurement noise in real-world systems, we conducted further experiments to evaluate robustness under more significant perturbations. Table 6 presents a comparison on all Hopper tasks with the deployment noise level increased to 10%. The results demonstrate that ROMBRL maintains its strong performance and continues to outperform strong baselines, highlighting its resilience against increased environmental uncertainty.
Table 6: Performance comparison on all Hopper tasks under a higher intensity of deployment noise (10% Gaussian noise). For each task, the “2nd-Best Algorithm” is the baseline that achieved the second-highest score under the original 5% noise condition (as detailed in Table 1). This comparison evaluates how previous top contenders perform under more significant perturbations. Results are averaged over 3 seeds.
Hopper Task ROMBRL (ours) 2nd-Best Algorithm
hopper-random 31.3 (0.7) 22.3 (1.0) [RAMBO]
hopper-medium 105.0 (0.0) 104.7 (0.6) [RORL]
hopper-medium-replay 108.0 (1.0) 99.7 (4.8) [MOBILE]
hopper-medium-expert 112.2 (0.8) 108.1 (7.9) [COMBO]
N.3 Training Curves
To provide a comprehensive view of the learning dynamics, this section presents the training curves for both the D4RL MuJoCo and the Tokamak Control benchmarks. Figure 6 illustrates the progression of evaluation scores on all D4RL MuJoCo tasks. Similarly, Figure 7 shows the learning curves for the three challenging Tokamak Control tasks. For both benchmarks, all policies are evaluated in their respective perturbed environments to assess deployment robustness throughout the training process.